OpenStudy (anonymous):

If f(-2)=4 and f'(-2)=-1, find the derivative of f(x)/x^2 at x=-2.

OpenStudy (anonymous):

\[g(x)=\frac{f(x)}{x^2}\] \[g'(x)=\frac{x^2f'(x)-2xf(x)}{x^4}\] by the quotient rule, fill in the numbers

OpenStudy (anonymous):

Thank you.